If the area of a right triangle is 9/16 sq. ft. And the height is 3/4 in., write an equation that relates the area to the base, b, and the height. Solve the equation to determine the base.
I tried .5(.75 x w) = .642
W= 1.708

Answer by Alan3354(69443)

If the area of a right triangle is 9/16 sq. ft. And the height is 3/4 in., write an equation that relates the area to the base, b, and the height. Solve the equation to determine the base.
-----------
Area = b*h/2
9/16 = (3/4)*b/2 = 3b/8
Multiply by 16 to eliminate fractions
9 = 6b
b = 3/2 feet
================
PS 9/16 is not .642, it's 0.5625